OpenStudy (anonymous):

A wooden garden tray with a base area of xy square inches is filled with soil up to a certain height. The tray contains (x2y + 10xy2 + 12xy) cubic inches of soil. The height of the tray is represented by the expression x squared y plus 10 x y squared plus 12 x y all over xy Which expression below is the correct simplified form of the height of the tray? a. 11xy + 12 b. x + 10y + 12 c. x + 10xy2 + 12xy d. x2y + 10xy2 + 12

OpenStudy (anonymous):

Since the volume of the garden tray is equal to the base area multiplied by the height (V=B*H), the final expression for volume is equal to xy*height. Height is equal to the volume of the tray divided by its base (H=V/B). To find the height, factor xy (base area) out of each term in the given expression for volume: V=(x^2)y + 10x(y^2) + 12xy=xy(x+10y+12), height=(x+10y+12).

OpenStudy (anonymous):

so it's b?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

your welcome :)